# The ubiquity of variable radio emission and spin-down rates in pulsars

### v1

@@ Abstract
Pulsars are often lauded for their (relative) rotational and radio emission stability over long time scales. However, long-term observing programmes are identifying an increasing number that deviate from this preconceived notion. Using Gaussian process regression and Bayesian inference techniques, we investigated the emission and rotational stability of 259 radio pulsars that have been monitored using Murriyang, the Parkes 64 m radio telescope, over the past three decades. We found 238 pulsars display significant variability in their spin-down rates, 52 of which also exhibit changes in profile shape. Including 23 known state-switching pulsars, this represents the largest catalogue of variable pulsars identified to date and indicates that these behaviours are ubiquitous among the wider population. The intensity of spin-down fluctuations positively scales with increasing pulsar spin-down rate, with only a marginal dependence on spin-frequency. This may have substantial implications for ongoing searches for gravitational waves in the ensemble timing of millisecond pulsars. We also discuss challenges in explaining the physical origins of quasi-periodic and transient profile/spin-down variations detected among a subset of our pulsars.
@@ 1 Introduction
The highly predictable, periodic nature of radio pulsars makes them excellent tools for studying a variety of astrophysical phenomena. These include measuring fluctuations in the electron column density and turbulence of the interstellar medium within our Galaxy (e.g. Petroff et al., 2013), testing theories of gravity (Freire & Wex, 2024), and the direct detection of nanohertz-frequency gravitational waves (Sazhin, 1978; Detweiler, 1979). Key to the success of such studies are the relative stability of both the pulsar emission mechanism and the rate at which pulsars spin down. While individual pulses from a given pulsar can vary wildly in intensity and shape, integrating over many rotations will typically return an average profile that is more or less constant from one observation to the next. Yet long-term monitoring programmes and observations with newer, more sensitive instruments have revealed a growing number of number of both mechanisms and cases that deviate from ideal clock assumptions.
@@
Mode or state switching between two or more discrete radio emission states was identified soon after the discovery of the first pulsars, manifesting as changes in the observed intensity, shape or position in pulse longitude of pulse profile components (see Backer 1970b, c; Lyne 1971). It is a broadband phenomenon, having been detected both at radio and X-ray wavelengths (Hermsen et al., 2013, 2017, 2018), and takes place over a variety of timescales. This can be as short as a one or several rotations of a neutron star, or slow secular variations that take place over many months or years. In some cases the emission appears to briefly switch off and back on again (nulls; Backer 1970a) or remain off for extended periods of time (intermittency; Kramer et al. 2006). Long-term monitoring experiments found that the timing of some pulsars displayed quasi-periodic processes with a range of apparent periods (e.g. Cordes & Greenstein 1981; Cordes & Downs 1985; Stairs et al. 2000). Both kinds of fluctuation in radio emission and spin-down rate were eventually linked, initially through observations of the intermittent pulsar PSR B1931$+$24 (Kramer et al., 2006), and later to increasingly large samples of pulsars that switch between two or more quasi-stable spin-down levels or undergo with transient spin-down events (Kramer et al., 2006; Hobbs et al., 2010; Lyne et al., 2010; Brook et al., 2014, 2016; Shaw et al., 2022). Other studies have revealed pulsars with variations in either spin-down or pulse shape (and even polarization) that appear uncorrelated with one another (Brook et al., 2016; Shaw et al., 2022; Basu et al., 2024), though this may be a selection effect imparted by limited telescope sensitivity and observing cadence.
@@
Exactly what mechanism drives these behaviours, and whether they originate from the interior dynamics or magnetospheric fluctuations of neutron stars, is presently unknown. Free precession of a non-axisymetric neutron star was once a popular hypothesis for explaining highly periodic changes in pulsar profile shapes and spin-down rates (Stairs et al., 2000). However, the discovery that such spin-down rate changes were linked to short-term emission variations has cast doubt on whether free precession is the true, or only, mechanism behind these behaviours (Lyne et al., 2010; Jones, 2012; Kerr et al., 2016; Stairs et al., 2019). Simultaneous emission switching at radio and X-ray wavelengths point toward a magnetospheric reconfiguration being responsible for (at least some) of the mode-switching behaviour (Hermsen et al., 2013). Emission and spin-down variations that seemingly coincided with glitches in several pulsars (e.g, Weltevrede et al. 2011; Keith et al. 2013) may point to crustal events such as starquakes that couple the internal dynamics of the neutron star to the magnetosphere (Akbal et al., 2015; Yuan et al., 2021). Indeed the 2016 glitch in the Vela pulsar (PSR J0835--4510/B0833--45) was associated with a short-lived quenching of its radio emission (Palfreyman et al., 2018), speculated to be due to a quake-induced magnetic disturbance (Bransgrove et al., 2020). External triggers such as interactions between pulsars and minor bodies are another popular hypothesis. These include interactions with individual in-falling asteroids (Brook et al., 2014), debris disks (Li, 2006; Shannon et al., 2013; Jennings et al., 2020b), and impacts from interstellar objects (Pham et al., 2024). In this case, the ionised remains of a wayward asteroid alter the plasma content of the pulsar magnetosphere, subsequently enhancing or attenuating the radio emission and altering the spin-down state. Many previous attempts to explain pulsar variability focussed on either individual objects or small samples with similar characteristics (e.g., Brook et al., 2014, 2016; Lower et al., 2023; Basu et al., 2024). Understanding whether one or more of these processes are responsible for these profile/spin-down variations requires population-level studies of a large ensemble of pulsars.
@@
Determining the prevalence of pulsar rotation/emission state changes across the population is vital to building a complete picture of the processes that contribute to the overall noise budget of a pulsar. Knowing how these behaviours scale as functions of spin frequency and spin-down rate is particularly important for pulsar timing array experiments that utilise an ensemble of millisecond pulsars to search for nanohertz-frequency gravitational waves. Even the most stable millisecond pulsars have been found to display profile shape variations and emission state switching that can hamper such experiments (Shannon et al., 2016; Brook et al., 2018; Goncharov et al., 2021; Miles et al., 2022; Nathan et al., 2023; Jennings et al., 2024). It may also bring us a step closer to determining the origins of unexpected time-dependent changes in the common noise process detected by pulsar timing array experiments (see Reardon et al. 2023).
@@ 2 Observations
The pulsar data analysed in this work were collected using Murriyang between 2007-2023 under the young pulsar timing (P574) programme (Weltevrede et al., 2010; Johnston et al., 2021). Pulsar observations were undertaken with the 10/50111Later re-tuned to cover the 40-cm band due to increased interference., 20-cm Multibeam, H-OH and Ultra-Wideband Low (UWL) receivers (Granet et al., 2001; Staveley-Smith et al., 1996; Granet et al., 2011; Hobbs et al., 2020) and saved to psrchive format archives (Hotan et al., 2004; van Straten et al., 2012). Details of the specific observing set-up, radio-frequency excision, and calibration steps can be found in Namkham et al. (2019); Parthasarathy et al. (2019); Lower et al. (2021); Johnston et al. (2021). Where possible, we also made use of extended timing datasets collected using legacy signal processors going back to the 1990s, which are available in the Parkes Observatory Pulsar Data Archive (Hobbs et al., 2011) [^2] https://data.csiro.au/domain/atnf.
@@ 3.1 Updated pulsar timing
A substantial fraction of the pulsars required updates to their timing models over those published in the ATNF pulsar catalogue (Manchester et al., 2005) [^3] https://www.atnf.csiro.au/research/pulsar/psrcat/. Phase coherent solutions were derived following the manual procedures outlined in Parthasarathy et al. (2019) and Lower et al. (2021), where pulsar positions, proper motions, spin frequency ($\nu$), spin-down ($\dot{\nu}$) and on occasion, second spin-frequency derivative ($\ddot{\nu}$), were re-fit using a combination of tempo2 and temponest (Hobbs et al., 2006; Lentati et al., 2014). The majority of pulsars that required inclusion of $\ddot{\nu}$ were those previously studied in Parthasarathy et al. (2020) and Lower et al. (2021). Glitch events were identified by eye as discontinuities or sharp features in the timing residuals. These were corrected for by fitting for the corresponding changes in $\nu$ and $\dot{\nu}$ using tempo2. Obtaining accurate measurements of the recovery parameters is challenging and imperfectly removed recovery signals will introduce unwanted artefacts in our spin-down analysis. A detailed analysis of the new glitch events and overall updates to the timing of these pulsars is beyond the scope of this work, so will be reported elsewhere. Once a coherent solution was obtained, we then applied pulse numbering to the pulse times of arrival (ToAs) to maintain an accurate record of every single rotation over our timing baseline. In Figure 1 we show the corresponding timing residuals for our final sample of 259 pulsars.
Figure 1: Timing residuals for all 259 pulsars after subtracting the best-fit timing model. Labels on the left-hand side of each panel indicates the pulsar J2000 name and the maximum-to-minimum range of the residuals in milliseconds.
@@
3.2 Modelling profile shape and spin-down variations
@@
We made use of the Matérn and white noise kernels implemented in scikit-learn (Pedregosa et al., 2011) for modelling the differences in profile shapes after subtracting a median template. This is the same approach that was used by Shaw et al. (2022), which is better at capturing short-duration profile shape changes than the squared-exponential kernel. For the Matérn kernel, we used a positive covariance exponent of $\mu=3/2$
@@
where $\sigma_{f}^{2}$ is the maximum allowed covariance, $|t-t^{\prime}|$ is the distance between two data points, and $\lambda$ is the kernel length scale. We computed the optimal kernel hyperparameters using a maximum-likelihood approach where the kernel length scale was allowed to vary between $30\,{\rm d}<\lambda<300\,{\rm d}$. An exception is PSR J1359--6038, which required two discrete length scales of $30\,{\rm d}<\lambda_{1}<600\,{\rm d}$ and $600\,{\rm d}<\lambda_{2}<3000\,{\rm d}$ to correctly capture the short and long-term profile variations displayed by this pulsar. In general, we found the scikit-learn implementations of the Matérn and white noise kernels provided more accurate representations of the underlying profile variations in our pulsars when compared to a similar variant implemented in the george Python package (Ambikasaran et al., 2015). The latter would often miss rapid changes in the profile shape.
Our approach to modelling the spin-down rates followed that of Shaw et al. (2022), where we used the gp_nudot.py script in PulsarPVC [^4] https://gitlab.com/benjaminshaw/pulsarpvc to fit a squared exponential kernel to the timing residuals for each pulsar. Spin-down timeseries were then generated by taking the second derivative of the resulting model of the timing residuals. As with the profile variability, this made use of the Gaussian process regression functionality of scikit-learn. Similar to the analyses of Brook et al. (2016) and Shaw et al. (2022), about half of our pulsars required the addition of a second squared exponential kernel to account for both long- and short-term variations. Several dozen pulsars in our sample displayed a sinusoidal signal with a 1-yr period in their resulting spin-down timeseries that resulted from small offsets from their assumed position on the sky. For some pulsars, this was corrected through updated timing model parameters obtained via tempo2 and TempoNest. Yet for others, a degeneracy between the annual sinusoid and quasi-periodic pulsar timing noise, combined with the limitations of the red power-law timing noise model in TempoNest, prevented us from making such positional corrections for a substantial number of pulsars. To overcome this, we also conducted a set of Gaussian process regression fits that included a sinusoidal kernel with a fixed 1-yr period. In order to determine which pulsars had residuals best described by a one or two kernel fit, or required an annual sinusoid to correct for positional offsets, we comparing the Bayesian information criterion (BIC) for each model. We first computed the BIC as
@@
where $k$ is the number of free parameters in the model, $N$ the number of data points, and $\mathcal{L}_{\max}$ is the maximum likelihood value for the data given the model. A smaller BIC indicates a better match between the model and the data. The statistical significance for which one model is preferred over the other can be inferred from the difference in BICs for each model
@@ 3.3 Identifying spin-down variations
where $\dot{\nu}_{\rm min}$ and $\dot{\nu}_{\rm max}$ refer to the minimum and maximum inferred spin-down derived from the Gaussian process regression, and $\sigma_{\dot{\nu},{\rm mean}}$ is the mean spin-down uncertainty. We used a threshold of $\mathcal{K}>1$ for defining when a pulsar displayed substantial spin-down variability.

### v2

@@ Abstract
Pulsars are often lauded for their (relative) rotational and radio emission stability over long time scales. However, long-term observing programmes are identifying an increasing number of pulsars that deviate from this preconceived notion. Using Gaussian process regression and Bayesian inference techniques, we investigated the emission and rotational stability of 259 radio pulsars that have been monitored using Murriyang, the Parkes 64 m radio telescope, over the past three decades. We found 238 pulsars that display significant variability in their spin-down rates, 52 of which also exhibit changes in profile shape. Including 23 known state-switching pulsars, this represents the largest catalogue of variable pulsars identified to date and indicates that these behaviours are ubiquitous among the wider population. The intensity of spin-down fluctuations positively scales with increasing pulsar spin-down rate, with only a marginal dependence on spin-frequency. This may have substantial implications for ongoing searches for gravitational waves in the ensemble timing of millisecond pulsars. We also discuss challenges in explaining the physical origins of quasi-periodic and transient profile/spin-down variations detected among a subset of our pulsars.
@@ 1 Introduction
The highly predictable, periodic nature of radio pulsars makes them excellent tools for studying a variety of astrophysical phenomena. These include measuring fluctuations in the electron column density and turbulence of the interstellar medium within our Galaxy (e.g. Petroff et al., 2013), testing theories of gravity (Freire & Wex, 2024), and the direct detection of nanohertz-frequency gravitational waves (Sazhin, 1978; Detweiler, 1979). The key to the success of such studies are the relative stability of both the pulsar emission mechanism and the rate at which pulsars spin down. While individual pulses from a given pulsar can vary wildly in intensity and shape (e.g. Johnston et al. 2001), integrating over many rotations will typically return an average profile that is more or less constant from one observation to the next. Yet long-term monitoring programmes and observations with newer, more sensitive instruments have revealed a growing number of cases that deviate from ideal clock assumptions.
@@
Mode or state switching between two or more discrete radio emission states was identified soon after the discovery of the first pulsars, manifesting as changes in the observed intensity, or shape or position in pulse longitude of the pulse profile components (see Backer 1970b, c; Lyne 1971). It is a broadband phenomenon, having been detected both at radio and X-ray wavelengths (Hermsen et al., 2013, 2017, 2018), and takes place over a variety of timescales. This can be as short as a one or several rotations of a neutron star, or slow secular variations that take place over many months or years. In some cases the emission appears to briefly switch off and back on again (nulls; Backer 1970a) or remain off for extended periods of time (intermittency; Kramer et al. 2006). Long-term monitoring experiments found that the timing of some pulsars displayed quasi-periodic processes with a range of apparent periods (e.g. Cordes & Greenstein 1981; Cordes & Downs 1985; Stairs et al. 2000). Both of the fluctuations in radio emission and spin-down rate were eventually linked, initially through observations of the intermittent pulsar PSR B1931$+$24 (Kramer et al., 2006), and later through increasingly large samples of pulsars that switch between two or more quasi-stable spin-down levels or undergo transient spin-down events (Kramer et al., 2006; Hobbs et al., 2010; Lyne et al., 2010; Brook et al., 2014, 2016; Shaw et al., 2022). Other studies have revealed pulsars with variations in either spin-down or pulse shape (and even polarization) that appear uncorrelated with one another (Brook et al., 2016; Shaw et al., 2022; Basu et al., 2024), though this may be a selection effect imparted by limited telescope sensitivity and observing cadence.
@@
Exactly what mechanism drives these behaviours, and whether they originate from the interior dynamics or magnetospheric fluctuations of neutron stars, is presently unknown. Free precession of a non-axisymmetric neutron star was once a popular hypothesis for explaining highly periodic changes in pulsar profile shapes and spin-down rates (Stairs et al., 2000). However, the discovery that such spin-down rate changes were linked to short-term emission variations has cast doubt on whether free precession is the true, or only, mechanism behind these behaviours (Lyne et al., 2010; Jones, 2012; Kerr et al., 2016; Stairs et al., 2019). Simultaneous emission switching at radio and X-ray wavelengths point toward a magnetospheric reconfiguration being responsible for at least some of the mode-switching behaviour (Hermsen et al., 2013). Emission and spin-down variations that seemingly coincided with glitches in several pulsars (e.g, Weltevrede et al. 2011; Keith et al. 2013) may point to crustal events such as starquakes that couple the internal dynamics of the neutron star to the magnetosphere (Akbal et al., 2015; Yuan et al., 2021). Indeed the 2016 glitch in the Vela pulsar (PSR J0835--4510/B0833--45) was associated with a short-lived quenching of its radio emission (Palfreyman et al., 2018), speculated to be due to a quake-induced magnetic disturbance (Bransgrove et al., 2020). External triggers such as interactions between pulsars and minor bodies are another popular hypothesis. These include interactions with individual in-falling asteroids (Brook et al., 2014), debris disks (Li, 2006; Shannon et al., 2013; Jennings et al., 2020b), and impacts from interstellar objects (Pham et al., 2024). In such cases, the ionised remains of a wayward asteroid alter the plasma content of the pulsar magnetosphere, subsequently enhancing or attenuating the radio emission and altering the spin-down state. Many previous attempts to explain pulsar variability focussed on either individual objects or small samples with similar characteristics (e.g., Brook et al., 2014, 2016; Lower et al., 2023; Basu et al., 2024). Understanding whether one or more of these processes are responsible for these profile/spin-down variations requires population-level studies of a large ensemble of pulsars.
@@
Determining the prevalence of pulsar rotation/emission state changes across the population is vital to building a complete picture of the processes that contribute to the overall noise budget of a pulsar. Knowing how these behaviours scale as functions of spin frequency and spin-down rate is particularly important for pulsar timing array experiments that utilise an ensemble of millisecond pulsars to search for nanohertz-frequency gravitational waves. Even the most stable millisecond pulsars have been found to display profile shape variations and emission state switching that can hamper such experiments (Shannon et al., 2016; Brook et al., 2018; Goncharov et al., 2021; Miles et al., 2022; Nathan et al., 2023; Jennings et al., 2024). It may also bring us a step closer to determining the origins of unexpected time-dependent changes in the common noise process seen by pulsar timing array experiments (see Reardon et al. 2023).
@@ 2 Observations
The pulsar data analysed in this work were collected using Murriyang between 2007-2023 under the ‘young’ pulsar timing (P574) programme (Weltevrede et al., 2010; Johnston et al., 2021). Our sample covers 259 isolated, non-recycled, rotation-powered pulsars with characteristic ages ranging from 1.6 kyr (PSR J1513--5908) to 0.4 Gyr (PSR J0820--4114). In Figure 1 we show the positions of our pulsar sample in period-period-derivative space alongside those of the broader ATNF Pulsar Catalogue (Manchester et al., 2005) [^1] https://www.atnf.csiro.au/research/pulsar/psrcat/. Note, there is a slight bias towards pulsars with high rotational kinetic energy loss rates due to the programme’s continued support of the Fermi mission (Smith et al., 2023). Pulsar observations were undertaken with the 10/50222Later re-tuned to cover the 40-cm band due to increased interference., 20-cm Multibeam, H-OH and Ultra-Wideband Low (UWL) receivers (Granet et al., 2001; Staveley-Smith et al., 1996; Granet et al., 2011; Hobbs et al., 2020) and saved to psrfits format archives (Hotan et al., 2004; van Straten et al., 2012). Details of the specific observing set-up, radio-frequency excision, and calibration steps can be found in Namkham et al. (2019); Parthasarathy et al. (2019); Lower et al. (2021); Johnston et al. (2021). Where possible, we also made use of extended timing datasets collected using legacy signal processors going back to the 1990s, which are available in the Parkes Observatory Pulsar Data Archive (Hobbs et al., 2011) [^3] https://data.csiro.au/domain/atnf. Pulse times of arrival (ToAs) were obtained following the prescription outlined previously in Parthasarathy et al. (2019), Lower et al. (2021) and Lower et al. (2023). We averaged the individual observations in time and frequency to form one-dimensional pulse profiles, which were then cross-correlated against a noise-free template using the Fourier-domain Monte Carlo method built into the pat tool in psrchive. Note that we employed observing-band specific templates for each pulsar.
Figure 1: A pulsar period ($P$) and period-derivative ($\dot{P}$) diagram, where the pulsars sample analysed in this work are highlighted by black circles. Light-grey circles indicate all known pulsars published in v2.5.1 of the ATNF Pulsar Catalogue.
@@ 3.1 Updated pulsar timing
A substantial fraction of the pulsars required updates to their timing models over those published in the ATNF Pulsar Catalogue. Phase coherent solutions were derived following the manual procedures outlined in Parthasarathy et al. (2019) and Lower et al. (2021), where pulsar positions, proper motions, spin frequency ($\nu$), spin-down ($\dot{\nu}$) and on occasion, second spin-frequency derivative ($\ddot{\nu}$), were re-fit using a combination of tempo2 and temponest (Hobbs et al., 2006; Lentati et al., 2014). The majority of pulsars that required inclusion of $\ddot{\nu}$ were those previously studied in Parthasarathy et al. (2020) and Lower et al. (2021). Glitch events were identified by eye as discontinuities or sharp features in the timing residuals. These were corrected for by fitting for the corresponding changes in $\nu$ and $\dot{\nu}$ using tempo2. Obtaining accurate measurements of the recovery parameters is challenging and imperfectly removed recovery signals will introduce unwanted artefacts in our spin-down analysis. A detailed analysis of the new glitch events and overall updates to the timing of these pulsars is beyond the scope of this work, and will be reported elsewhere. Once a coherent solution was obtained, we then applied pulse numbering to the ToAs to maintain an accurate record of every single rotation over our timing baseline. In Figure 2 we show the corresponding timing residuals for our final sample of 259 pulsars.
Figure 2: Timing residuals for all 259 pulsars after subtracting the best-fit timing model. Labels on the left-hand side of each panel indicates the pulsar J2000 name and the maximum-to-minimum range of the residuals in milliseconds. The downward pointing arrows indicate the epochs of detected glitches.
@@
3.2 Profile variability maps
@@
We made use of the Matérn and white noise kernels implemented in scikit-learn (Pedregosa et al., 2011) for modelling the differences in profile shapes after subtracting a median template. This is the same kernel that was used by Shaw et al. (2022), which is better at capturing short-duration profile shape changes than the squared-exponential kernel. For the Matérn kernel, we used a positive covariance exponent of $\mu=3/2$
@@
where $\sigma_{f}^{2}$ is the maximum allowed covariance, $|t-t^{\prime}|$ is the time separating two data points, and $\lambda$ is the kernel length scale. We computed the optimal kernel hyperparameters using a maximum-likelihood approach where the kernel length scale was allowed to vary between $30\,{\rm d}<\lambda<300\,{\rm d}$. An exception is PSR J1359--6038, which required the use of two separate kernels with respective length scales of $30\,{\rm d}<\lambda_{1}<600\,{\rm d}$ and $600\,{\rm d}<\lambda_{2}<3000\,{\rm d}$ to correctly capture the short and long-term profile variations displayed by this pulsar. In general, we found the scikit-learn implementations of the Matérn and white noise kernels provided more accurate representations of the underlying profile variations in our pulsars when compared to a similar variant implemented in the george Python package (Ambikasaran et al., 2015). The latter would often miss rapid changes in the profile shape.
3.3 Modelling spin-down rate variations
Our approach to modelling the spin-down rates followed that of Shaw et al. (2022), where we used the gp_nudot.py script in PulsarPVC [^4] https://gitlab.com/benjaminshaw/pulsarpvc to fit a squared exponential kernel to the timing residuals for each pulsar. Spin-down timeseries were then generated by taking the second derivative of the resulting model of the timing residuals. As with the profile variability, this made use of the Gaussian process regression functionality of scikit-learn. Similar to the analyses of Brook et al. (2016) and Shaw et al. (2022), about half of our pulsars required the addition of a second squared exponential kernel to account for both long- and short-term variations. Several dozen pulsars in our sample displayed a sinusoidal signal with a 1-yr period in their resulting spin-down timeseries that resulted from small offsets from their assumed position on the sky. For some pulsars, this was corrected through updated timing model parameters obtained via tempo2 and TempoNest. Yet for others, a degeneracy between the annual sinusoid and quasi-periodic pulsar timing noise, combined with the limitations of the red power-law timing noise model in TempoNest [^5] Modelling of quasi-periodic and non-stationary timing behaviours as a red power-law process can result in biased measurements of the pulsar properties (cf. Keith & Niţu 2023)., prevented us from making such positional corrections for a substantial number of pulsars. To overcome this, we also conducted a set of Gaussian process regression fits that included a sinusoidal kernel with a fixed 1-yr period. In order to determine which pulsars had residuals best described by a one-kernel or two-kernel fit, or required an annual sinusoid to correct for positional offsets, we compared the Bayesian information criterion (BIC) for each model. We first computed the BIC as
@@ 3.3 Modelling spin-down rate variations
where $k$ is the number of free parameters in the model, $N$ is the number of data points, and $\mathcal{L}_{\max}$ is the maximum likelihood value for the data given the model. A smaller BIC indicates a better match between the model and the data. The statistical significance for which one model is preferred over the other can be inferred from the difference in BICs for each model
@@
where $\dot{\nu}_{\rm min}$ and $\dot{\nu}_{\rm max}$ refer to the minimum and maximum inferred spin-down derived from the Gaussian process regression, and $\sigma_{\dot{\nu},{\rm mean}}$ is the mean spin-down uncertainty computed via equations 9 and 10 of Brook et al. (2016). We used a threshold of $\mathcal{K}>1$ for defining when a pulsar displayed substantial spin-down variability. To test the level of correlation between any observed variations in profile shape and spin-down rate, we computed the Spearman rank correlation coefficient between each phase bin of the profile variability map and $\dot{\nu}$ for time-lags ranging between $\pm$ half the length of the per-pulsar observing span (Brook et al., 2016; Shaw et al., 2022; Lower et al., 2023).